OpenStudy (anonymous):

Can you solve the following inequality? x minus 4 over x minus 2....this fraction is less than or equal to negative 3 x−4 / x−2 ≤−3

OpenStudy (anonymous):

=(x-4)≤-3(x-2) =x-4≤-3x+6 =4x≤10 =>x=10/4=5/2

OpenStudy (anonymous):